OpenStudy (anonymous):

-(z+7)+(3z+1)=-5(z+1) the solution to the set is z=

OpenStudy (anonymous):

@ConDawg @GoldPhenoix @whpalmer4

OpenStudy (whpalmer4):

\[-(z+7)+(3z+1) = -5(z+1)\]apply distributive property \[-z-7 + 3z + 1 = -5z -5\]collect like terms \[2z-6 = -5z -5\]solve for \(z\)

OpenStudy (anonymous):

@whpalmer4 thats what i got as far as the 2z-6 but i dont understand anything else

OpenStudy (whpalmer4):

Okay, why don't you add 5z to both sides, what do you get?

OpenStudy (anonymous):

what 1z

OpenStudy (anonymous):

i mean 3z

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

huh? write out the equation, and add +5z on each side of the equals sign. what do you get?

OpenStudy (anonymous):

2z-6=5z-5 5z 5z 3z-6=0-5

OpenStudy (anonymous):

@whpalmer4 idk?

OpenStudy (anonymous):

@whpalmer4 Did most of the work for you here.

OpenStudy (anonymous):

\[2z-6=-5z-5\]

OpenStudy (anonymous):

You want you variables on one side and constants on the other. You can do that right?

OpenStudy (anonymous):

so is the solution 2z-6?

OpenStudy (anonymous):

No. Right now we have 2z-6=-5z-5. We need to know the value of just z. 2z-6=-5z-5 doesn't give us that value of z But we can change the equation 2z-6=-5z-5 to give us the value of z All you need to do is get the z's on one side of the equal sign and constants on the other

OpenStudy (anonymous):

or is it 0.1428571429

OpenStudy (anonymous):

well lets see

OpenStudy (anonymous):

2z-6=-5z-5 2z=-5z+1 7z=1 z= 1/7 1/7 = 0.1428571429 But it's better to use 1/7

OpenStudy (anonymous):

y is that @ConDawg

OpenStudy (anonymous):

Why is it better to use 1/7?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Because it's a nice fraction, and is just better to look at then 0.1428571429

OpenStudy (anonymous):

ok now i c @ConDawg

OpenStudy (anonymous):

:)

OpenStudy (whpalmer4):

sorry, had some technical difficulties. what I was trying to show you, @Avon, was \[2z-6=-5z-5\]We add \(5z\) to each side \[2z-6+5z = -5z - 5 + 5z\]now collect like terms \[7z - 6 = -5\]Now add 6 to each side \[7z - 6 + 6 = -5 + 6\]collect like terms \[7z = 1\]now divide each side by 7 to get \(z\) all alone \[7z/7 = 1/7\]\[z=1/7\]

OpenStudy (whpalmer4):

as long as we do the same thing to both sides of the equation, it remains true. if you and your friend both have $10, you both have the same amount of money. If someone comes along and gives each of you another $20 bill, you both have the same amount of money, even though the amount is now different than what it was originally.